Question 17

  1. A researcher wishes to test the claim that the average age of lifeguards in Ocean City is 24 years. The researcher selects a sample of 29 guards and finds the mean age to be 24.6 years with a standard deviation of 1.85 years. The level of significance is 10%. Assume the population is normally distributed.

    A. Find the critical value(s).

    B. Find the rejection region(s).

    Label your answers as A and B.

    You must state complete answers for both parts.
Answer #1

Solution:

This is two tailed test.

Degrees of freedom = n - 1 = 28

Level of significance = = 0.10

A.

Critical value = t/2,df

= t0.05,28

= 1.701

B.

Rejection region =

t < -1.701 or t > 1.701